cody's bank contains $4.30 in nickels, dimes and quarters there are 38 coins in all, with 6 more nickels than quarters how many dimes are there
Answer by stanbon(75887)

cody's bank contains $4.30 in nickels, dimes and quarters there are 38 coins in all, with 6 more nickels than quarters how many dimes are there
Quantity Eq: n + d + q = 38 coins
Quantity Eq: n = q + 6
Value Eq::: 5n +10e+25q = 430 cents
------------------------------------------
Rearrange
d + q + q+6 = 38
5(q+6)+10d + 25q = 430
------
Simplify:
d + 2q = 32
10d+30q = 400
-----
Modify for Elimination:
d + 2q = 32
d + 3q = 40
---
Subtract to solve for "q";
q = 8 (# of quarters)
---
Solve for "d":
d + 2q = 32
d + 16 = 32
d = 16 (# of dimes)
=====================
Cheers,
Stan H.